A time study analyst observes a worker over a period of time and calculates that the worker required 60.0 minutes per unit produced. The worker was rated by the analyst at a 120 percent performance rating, and an 15 percent allowance for personal time and fatigue applies.

(1) What is the standard time for this task?

(2) If a worker managed to produce 10 units in an eight-hour day, what would their day's pay be if the base rate is $19 per hour, and the premium payment system paid on a 100 percent basis?
Answer #1

Normal time = Observed time * PR

Where observed time = 60 min

PR = 120 %

NT = 60 * 120% = 72 minutes

Standard time = Normal time + Allowances * Normal time

Standard time = 72 + 15%*72 = 82.8 minutes

B. Pay:

Total time = 82.8 * 10 = 828 minutes or 828/60 = 13.8 hours

Hourly rate = $19

Pay = 13.8 * 19 = $ 262.2